OpenStudy (anonymous):

Austin wants to sketch the graph of a circle represented by the given equation. (x - 2)2 + (y + 5)2 = 36 Using complete sentences, describe the steps used to sketch the circle on a coordinate grid.

OpenStudy (mathstudent55):

The equation of a circle is \( (x - h)^2 + (y - k)^2 = r^2 \) where the center of the circle is (h, k) and the radius is r.

OpenStudy (anonymous):

Yeah but this question is confusing.

OpenStudy (mathstudent55):

I don't think it's that confusing. All you need to do is comare the given equation with the standard equation I wrote above and compare the numbers.

OpenStudy (anonymous):

I'm doing that, I just don't know how to write it in the box.

OpenStudy (anonymous):

...

OpenStudy (mathstudent55):

\( \large{ (x - 2)^2 + (y + 5)^2 = 36 }\) \( \large{ (x - \color{blue}2^2 + (y - (\color{red}{- 5}))^2 = \color{green}6^2 }\) \( \large{(x - \color{blue}h)^2 + (y - \color{red}k)^2 = \color{green}r^2 }\)

OpenStudy (anonymous):

Thank you.(:

OpenStudy (mathstudent55):

That means for your problem, the center is at (2, -5) and the radius is 6.

OpenStudy (anonymous):

I figured that one out, thank you for your help.(:

OpenStudy (mathstudent55):

You're welcome.